### v2

# LLMs Judging LLMs: A Simplex Perspective

Abstract
Given the challenge of automatically evaluating free‐form outputs from large language models (LLMs), an increasingly common solution is to use LLMs themselves as the judging mechanism, without any gold-standard scores. Implicitly, this practice accounts for only sampling variability (aleatoric uncertainty) and ignores uncertainty about judge quality (epistemic uncertainty). While this is justified if judges are perfectly accurate, it is unclear when such an approach is theoretically valid and practically robust. We study these questions for the task of ranking LLM candidates from a novel geometric perspective: for $M$-level scoring systems, both LLM judges and candidates can be represented as points on an (M1)(M-1)-dimensional probability simplex, where geometric concepts (e.g., triangle areas) correspond to key ranking concepts. This perspective yields intuitive theoretical conditions and visual proofs for when rankings are identifiable; for instance, we provide a formal basis for the “folk wisdom” that LLM judges are more effective for two-level scoring ($M=2$) than multi-level scoring ($M>2$). Leveraging the simplex, we design geometric Bayesian priors that encode epistemic uncertainty about judge quality and vary the priors to conduct sensitivity analyses. Experiments on LLM benchmarks show that rankings based solely on LLM judges are robust in many but not all datasets, underscoring both their widespread success and the need for caution. Our Bayesian method achieves substantially higher coverage rates than existing procedures, highlighting the importance of modeling epistemic uncertainty.
LLMs Judging LLMs: A Simplex Perspective
Patrick Vossler1∗          Fan Xia1∗          Yifan Mai2          Adarsh Subbaswamy3          Jean Feng1
1University of California, San Francisco          2Stanford University          3University of Maryland, Baltimore
11footnotetext: Equal contribution
1 Introduction
Scalable benchmarking of large language models (LLMs) is increasingly critical given the rapid proliferation of models, model updates, and new benchmark datasets. While multiple choice or numerical answers can be verified algorithmically, many benchmark tasks allow free-form text responses that are more difficult to verify, such as clinical reasoning expressed as natural language to mathematical proofs involving multi-part LaTeX formulas. While the current gold standard of consensus voting by multiple human experts is effective, it is often prohibitively expensive and difficult to scale.
In response to these challenges, recent work has proposed using LLMs themselves as judges (Figure 1) (Zheng2023-wp). We refer to evaluator LLMs as “judges” and evaluated LLMs as “candidates.” This now-common practice effectively treats LLM judges as perfectly accurate, ignoring any uncertainty about their quality and acknowledging only sampling variability (e.g., through bootstrapping or confidence intervals). Borrowing terms from epistemology, the first source of uncertainty is systematic or epistemic; due to the lack of gold-standard data, one is uncertain about and cannot determine the true quality of the judges. The second source of uncertainty is stochastic or aleatoric; due to sampling variability, there is noise in the returned judge scores but this type of uncertainty can be reduced by increasing sample size. The goal of this work is to understand how epistemic uncertainty impacts our ability to rank LLM candidates. We do so by answering two questions: (i) from a theoretical standpoint, what are sufficient conditions for an LLM-as-a-judge pipeline to identify the true rankings and (ii) from a methodological standpoint, how can we quantify our confidence in estimated rankings when we do not know if these theoretical conditions hold in practice?
Figure 1: LLM judge workflow: For each benchmark question, LLM judges score each candidate’s answer according to a rubric. Candidates are ranked based on their judge-assigned scores. Shaded boxes indicate cases where the same LLM serves as both candidate and judge (self-judging).
To this end, we show how a simple geometric perspective can illuminate the role of epistemic uncertainty and guide both theoretical and methodological development. The key idea is that we can simultaneously visualize judge quality and candidate score distributions on a probability simplex, where geometric concepts (e.g., triangle areas) correspond to key ranking concepts. Epistemic uncertainty manifests as uncertainty in the location of judge points, while aleatoric uncertainty manifests as stochasticity of candidate points. This representation allows us to reason about how the relative positions of these points affect identifiability and estimation of the true rankings.
Leveraging this geometric framework, we find that constancy of LLM judge behavior is sufficient for ranking identifiability in two-level scoring systems. However, this sufficiency breaks down when ranking across three or more levels, uncovering a surprising phase transition. These theoretical results motivate the development of a Bayesian framework for estimating ranks, where we construct minimally-informative geometric priors that encode our prior beliefs about judge quality. We then conduct sensitivity analyses of estimated rankings by varying the strength of the Bayesian priors, which corresponds to varying the level of epistemic uncertainty.
Applying our framework to five LLM benchmark datasets, our sensitivity analyses reveal that LLM rankings are robust in some cases (e.g., GPQA and SummEval) but not others (e.g., Omni-MATH). These results explain the practical success of LLM judges, but also emphasize the importance of their careful, measured use. Furthermore, the Bayesian framework achieves substantially higher coverage of the true rankings than existing methods, as it captures both aleatoric and epistemic uncertainty whereas standard approaches capture only the former. Together, these results underscore the important role of epistemic uncertainty when using LLM-as-a-judge pipelines to rank candidates and the need to incorporate this source of uncertainty to ensure statistically sound conclusions.
2 Related work
LLM-as-a-judge and their implicit assumptions. While LLM judges have emerged as a scalable alternative to traditional metrics that require reference outputs (Papineni2001-dx; Lin2004-bz; Wang2023-hy; Chiang2023-we; Li2024-ee; Gu2024-im), current approaches may not fully capture all sources of uncertainty in their rankings. Despite documented biases in LLM judges (Zheng2023-wp; Koo2024-gd; Wei2025-mp), proposed mitigations—score averaging, juries, judge debates, rubrics (Liang2023-ow; Verga2024-vf; Kalra_undated-rn; Chan2024-dj; Databricks2024-rb; Lee2024-qb)—implicitly assume that judges are perfect or that their errors cancel out in expectation, ignoring uncertainty about judge quality. While Guerdan2025-to discussed the limitations of LLM judges when there is no true agreed-upon rating scale, we prove that even with agreed-upon scales, fundamental identifiability limits exist that no averaging or ensemble method can overcome.
Uncertainty quantification without ground truth. Existing uncertainty quantification methods cannot address the fundamental challenge of epistemic uncertainty about judge quality. Bootstrapping (Goldstein1996-mj; Xie2009-wv) only quantifies sampling variation, leading to undercoverage when assumptions fail. Methods like prediction-powered inference (Chatzi2024-cy; Angelopoulos2023-tv), conformal inference (Jung2025-lt), or consensus models (Raykar2010-dd; Welinder2010-rc) can integrate LLM judging if some subset of gold-standard labels are available for each new model and benchmark, limiting scalability. Bradley-Terry models (Bradley1952-kj; Rao1967-fv; Herbrich2006-jx; Ameli2025-fg), widely used in rating systems like Elo and Chatbot Arena, assume strong stochastic transitivity, restricting their applicability to settings where pairwise comparisons uniquely determine a global ranking.
Imperfect reference standards. Using LLMs as imperfect judges parallels evaluation of medical diagnostics with imperfect reference standards (Reitsma2009-rn; Umemneku_Chikere2019-kt; Sun2025-si; Sun2024-sj). Our results expand on ideas used in this field, beyond the classical setting of evaluating binary diagnostic tests to the case of multi-level ratings, which are commonly used for LLM evaluation. The geometric arguments substantially extend (Fienberg1970-dt; Black2002-wq; Jones2010-qa; Duan2020-lg), which focused on identifying diagnostic test performance rather than ranking and only considered the case with binary diagnostic tests.
3 A geometric perspective
We study the LLM judge pipeline shown in Figure 1, where $J$ LLM judges score answers by $K$ candidate LLMs to questions from a benchmark task. Formally, denote the spaces of questions and answers as $\mathcal{Q}$ and $\mathcal{A}$, respectively. A benchmark task is defined by a distribution over questions $Q\sim P_{\mathcal{Q}}$ and a true scoring function $s^{*}:\mathcal{Q}\times\mathcal{A}\to\{1,\ldots,M\}$. Each candidate $k$ is a (potentially stochastic) function $f_{k}:\mathcal{Q}\to\mathcal{A}$ that generates answers, with true score $S^{*}_{k}=s^{*}(Q,f_{k}(Q))$ for question $Q$. Candidate $k$’s true score distribution is denoted $\pi_{k}=(\pi_{k,1},\ldots,\pi_{k,M})$, where $\pi_{k,m}=\Pr(S^{*}_{k}=m)$ is the frequency with which the candidate achieves true score $m$. The goal is to recover the true ranking of candidates with respect to their expected true scores $\mathbb{E}[S^{*}_{k}]=\sum_{m=1}^{M}m\cdot\pi_{k,m}$.
In the setting where one does not observe the true scores or the true scoring function, we only have access to the LLM judge pipeline to obtain judge-assigned scores. Each judge $j=1,\cdots,J$ is represented as a function $\hat{s}_{j}:\mathcal{Q}\times\mathcal{A}\to\{1,\ldots,M\}$. So given question $Q$ and candidate $k$’s answer, the judge assigns score $\hat{S}_{k}^{(j)}=\hat{s}_{j}(Q,f_{k}(Q))$. Given infinite questions drawn from the benchmark task, the distribution of judge-assigned scores for candidate $k$ is $\gamma^{(j)}_{k}=(\gamma^{(j)}_{k,1},\ldots,\gamma^{(j)}_{k,M})$, where $\gamma^{(j)}_{k,m}=\Pr(\hat{S}_{k}^{(j)}=m)$ is the frequency that judge $j$ assigns score $m$ to the candidate’s answers.
Figure 2: The true prevalence of the true scores corresponds to barycentric coordinates; (a) and (b) highlight prevalence $\pi_{1}$ in 2- and 3-level scoring systems, respectively. A candidate’s expected score corresponds to the height of vertical projection in the augmented space, as illustrated in (c).
Using simple geometric arguments, we can relate the true score distribution $\pi_{k}$ to the observed score distribution $\gamma_{k}^{(j)}$ using judge $j$’s confusion matrix for candidate $k$. In particular, let the $m$th column of the judge’s confusion matrix for candidate $k$ represent the distribution of assigned scores for answers with true score $m$, i.e.,
$\theta^{(j)}_{m,k}=\big(\Pr(\hat{S}_{k}^{(j)}=1|S^{*}_{k}=m),\ldots,\Pr(\hat{S}_{k}^{(j)}=M|S^{*}_{k}=m)\big).$
Each of the judge’s confusion matrix columns $\theta^{(j)}_{m,k}$ for $m=1,\cdots,M$ lie on the (M1)(M-1)-dimensional probability simplex and, critically, the candidate’s judge-assigned score distribution $\gamma_{k}^{(j)}$ is the convex combination of these columns with respect to true score distribution, i.e., $\gamma^{(j)}_{k}=\sum_{m=1}^{M}\pi_{k,m}\theta^{(j)}_{m,k}$. As examples, Figure 2a shows how the judge and candidate in a 2-level scoring system all fall on a line segment and Figure 2b shows how these points in a 3-level scoring system all fall within a triangle. Using this geometric perspective, we can then establish equivalences between geometric concepts to key ranking concepts:
1. True score distributions correspond to barycentric coordinates. $\pi_{k}$ corresponds precisely to the barycentric coordinates for the weighted centroid $\gamma^{(j)}_{k}$ relative to judge vertices $\theta^{(j)}_{m,k}$, which are also known as areal coordinates because they can be expressed as ratios of simplex sub-areas. For instance, for $M=3$, $\pi_{k,1}$ equals the area of the subtriangle from the weighted centroid to the 2nd and 3rd vertices divided by the area of the triangle defined by all vertices (Figure 2b), i.e.,
$\pi_{k,1}=\frac{\text{Area of }\triangle_{(\gamma_{k},\theta_{2},\theta_{3})}}{\text{Area of }\triangle_{(\theta_{1},\theta_{2},\theta_{3})}},$ (1)
with $\pi_{k,2}$ and $\pi_{k,3}$ defined analogously. This geometric characterization extends to any $M$.
2. Expected scores map to height in augmented space. If we augment the probability simplex with an additional axis representing score values and lift each judge vertex $\theta^{(j)}_{m,k}$ to height $m$, the vertices form an augmented simplex (Figure 2c). Then, as proved in the Appendix, candidate $k$’s expected score $\mathbb{E}[S^{*}_{k}]$ corresponds precisely to the height of its vertical projection onto this augmented simplex.
3. Epistemic uncertainty manifests as unknown positions of judge vertices. When multiple configurations of judge vertices are compatible with the observed set of candidate points, the location of the judge vertices cannot be determined. Thus epistemic uncertainty of judge quality is represented by uncertainty in the location of judge vertices.
4. Aleatoric uncertainty manifests as stochasticity of candidate points. Aleatoric uncertainty is sampling variability, which corresponds to us observing the empirical judge-assigned score distributions $\hat{\gamma}_{k,n}^{(j)}$ for a benchmark dataset with $n$ questions, rather than the expected judge-assigned score distribution $\gamma_{k}^{(j)}$.
4 Theoretical limits of ranking identifiability
Figure 3: Visualization of judge assumptions for the 2-level scoring setting. We suppose there are two LLMs (1=blue, 2=orange), where both are candidates (with score distributions $\gamma_{1}$ and $\gamma_{2}$) and LLM 1 is a judge. (a) Perfect Judge assumes judge vertices $(\theta_{1},\theta_{2})$ are at the extremes of the 1-dimensional probability simplex. (b) Strong Constancy assumes the vertex positions for the judge are same across all candidates. (c) Moderate Constancy assumes the vertex positions for the judge only differ when self-judging (indicated by the red shifted vertices).
Leveraging this geometric framework, we now investigate when true rankings can or cannot be recovered from LLM judge scores without ground truth labels. To understand the theoretical limits, we assume access to infinite data samples, resulting in zero aleatoric uncertainty and leaving only epistemic uncertainty.
In the most general case with zero constraints on the LLM judge behaviors, our geometric visualization will have $J\times K\times M$ points representing the performance of the $J$ judges across the $K$ candidates as well as $J\times K$ points representing the score distributions assigned by each judge to each candidate. The true rankings are generally nonidentifiable in this setting, so we must impose structure on judge behavior. Thus we consider assumptions, that when imposed, may sufficiently reduce epistemic uncertainty to make the true rankings identifiable.
The strongest assumption is to assume the judges are perfect. The confusion matrices collapse to the identity matrix, placing all judge vertices at the simplex corners (Figure 3a). If true, this assumption would obviously make the true rankings identifiable for any value of $J$, $M$, and/or $K$. However, this is likely unrealistic.
A weaker and somewhat more realistic assumption is that the judge’s performance is perfectly consistent across candidates (Strong Constancy). Mathematically, this corresponds to the confusion matrix being the same across candidates, i.e., the vertices for each judge $j$ are shared across candidates (Figure 3b).
Assumption 1 (Strong Constancy).
The confusion matrix of judge $j$ is identical across all $K$ candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for $k=1,\cdots,K$.
A weaker assumption that allows for self-preference bias (Koo2024-gd) is to suppose that the LLM judge is perfectly consistent as long as it is not judging itself (Moderate Constancy). This can be visualized by a separate set of vertices when the LLM is judging itself but a shared set of vertices otherwise (Figure 3c).
Assumption 2 (Moderate Constancy).
Judge $\hat{s}_{j}$ has an identical confusion matrix for all non-self candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for all $k\neq j$. [^1] For simplicity, the theoretical results studies the case where self-judging means the same LLM judges itself. In practice, we exclude judges from judging their model family.
While one may consider even weaker assumptions, we will study the implications of these two constancy conditions on ranking identifiability, as even here the results are nuanced. These theoretical results will then motivate our framework in the next section for estimating ranks and quantifying uncertainty.
4.1 2-level scoring systems
Let us consider the simplest setting: a 2-level scoring system with a single judge that satisfies Strong Constancy. In this setting, all candidates must lie on the line segment between the two judge vertices $\theta_{1}$ and $\theta_{2}$. Figure 2 provides visual proof that the true rankings, which is equivalent to ranking prevalences of score $m=2$, is given by the relative positions of the candidates along the line. This is true even though we do not know the exact positions of the judge vertices and even if the LLM judges are far from perfect. In fact, our only requirement is that the judges are not adversarial (to avoid the so-called “label-flipping problem” (Sun2025-si)):
Assumption 3.
Judge $j$’s probability of assigning the lowest score when the true score is equal to $m$ decreases with respect to $m$.
The story under Moderate Constancy is similar. Using similar geometric reasoning (see Appendix), we find that the only difference is that at least two judges and four candidates are needed to obtain enough information to recover the true ranking.
Theorem 1.
For 2-level scoring, any of the following conditions are sufficient for candidate rankings to be identifiable from the distribution of judge-assigned scores: (i) There is $J=1$ judge and it satisfies Assumptions 4 and 3. (ii) There are $K\geq 4$ candidates, $J\geq 2$ judges, and the judges satisfy Assumptions 5 and 3.
Remark. For completeness, we present in the Appendix necessary and sufficient conditions for recovering rankings in two-level systems. The geometry is complex, so we defer the details to the appendix.
4.2 3+-level scoring systems
The story is more nuanced for scoring systems with 3 or more levels. Below, we present geometric arguments for 3-level scoring systems, but the same arguments extend to more levels. To build intuition, we first discuss the tasks of (i) ranking the prevalence of each score and then (ii) ranking the expected scores.
(i) Ranking prevalences. Without loss of generality, suppose we wanted to rank the prevalence of true score $m=1$. As mentioned in Section 3, this is equivalent to ranking the barycentric coordinates $\pi_{k,1}$ across candidates $k=1,\cdots,K$, which are defined as the area ratios (1). Under strong constancy, the denominators of these ratios are the same for all $k$, so ranking $\pi_{k,1}$ reduces to ranking the areas of the subtriangles $\triangle_{(\gamma_{k},\theta_{2},\theta_{3})}$ Nevertheless, we find that the rankings are non-identifiable in general, even under strong constancy. Figure 4 provides a visual proof: panels (a) and (b) show two equally valid judge configurations ($\{\theta_{1},\theta_{2},\theta_{3}\}$ versus $\{\theta^{\prime}_{1},\theta^{\prime}_{2},\theta^{\prime}_{3}\}$) that can explain candidate score distributions $\gamma_{k=1}$ and $\gamma_{k=2}$, as both judge configurations envelop the candidate points. However, the areas flip between configurations. In panel (a), subtriangle $\triangle_{(\gamma_{1},\theta_{2},\theta_{3})}$ has greater area than $\triangle_{(\gamma_{2},\theta_{2},\theta_{3})}$, while in panel (b) this relationship reverses. Thus the two panels show that even under strong constancy, the rankings of the score prevalences cannot be recovered if there is no prior knowledge about judge quality.
Figure 4: Non-identifiability in 3-level scoring. Same candidate positions (green/red) explained by different judge configurations yield opposite prevalence rankings: $\pi_{1,1}>\pi_{2,1}$ in (a) vs. $\pi_{1,1}<\pi_{2,1}$ in (b).
(ii) Ranking expected scores. Given the nonidentifiability of prevalence rankings, it is unsurprising that the ranking of true candidate scores is also nonidentifiable when gold standard labels are unavailable. To see this more directly, notice that the expected scores of the candidates, which correspond to the relative heights of each candidate’s vertical projection onto the augmented simplex (Figure 2c), varies based on the location of the judge vertices. If a judge’s vertices are shifted enough, the height rankings between candidates can shift. This argument applies, under either constancy assumption and for any number of judges.
Together, these results illustrate that epistemic uncertainty can significantly affect recoverability of the true rankings, even when aleatoric uncertainty is removed.
Theorem 2.
Consider the 3+-level scoring setting with all judges either satisfying Assumptions 4 and 3 or Assumptions 5 and 3. Given only the distribution of judge-assigned scores, there exist candidates whose prevalence and/or expected-score rankings cannot be identified.
4.3 Recoverability is dataset-specific
The prior theoretical results show that recoverability depends not only on judge quality assumptions but also the problem setting. In particular, Theorems 1 and 2 uncover a dependence on the number of scoring levels, supporting the “folk wisdom” that LLM judges tend to be more effective for 2-level scoring systems and less effective for 3+levels (Shankar2024-an; Husain2025-js). Leveraging the geometric perspective further, we find that recoverability is, in fact, dataset-specific, as illustrated in the following two examples. This means we need a way to rigorously quantify how sensitive the estimated rankings for a given dataset are to varying levels of epistemic uncertainty, which we address using Bayesian inference in the next section.
Example 1: Dataset requiring stronger priors. Consider again the Figure 4 example and the task of ranking prevalence of score $m=1$. Intuitively, the judge configuration in panel (a) may seem more probable than panel (b) if we believe that LLM judges can distinguish scores 2 and 3 moderately well. To formalize this prior belief, we distill it using geometric arguments. In particular, ranking $\pi_{k,1}$ is equivalent to ranking areas of subtriangles $\triangle{\gamma_{k},\theta_{2},\theta_{3}}$. Because all the subtriangles have the same line segment $\overline{\theta_{2}\theta_{3}}$ as the base, ranking areas is equivalent to ranking distances from each candidate $\gamma_{k}$ to the line $\overrightarrow{\theta_{2}\theta_{3}}$. Because ranking distances only require knowing the slope of this line, it is sufficient to state our prior belief about judge quality in terms of slopes. This is something we generally have a good prior about. In contrast, it is generally more difficult to specify one’s prior about the exact location of the judge vertices, since LLMs are often miscalibrated (Wang2024-gp).
Example 2: Dataset allowing weaker priors. Consider a candidate receiving mostly 1’s versus another receiving mostly 3’s from an LLM judge (see visualization in Appendix). When score distributions are so different, the ranking between two candidates may be quite evident, even if we have limited prior knowledge on the performance of the LLM judges and even if the constancy assumptions do not hold. However, we need a method to quantify when constancy violations flip rankings, as discussed in the next section.
5 Bayesian framework
Our theoretical findings and geometric framework have shown us that how much epistemic uncertainty impacts our ability to rank LLM candidates is dataset-specific. To quantify this more formally, we have designed a Bayesian framework that encodes epistemic uncertainty in terms of two priors: (i) how strongly the constancy assumption holds, as represented through random effects (Section 5.2), and (ii) prior beliefs about judge discriminative ability, as represented by slopes between judge vertices (Section 5.3). We can then conduct sensitivity analyses that assess how robust the estimated rankings are to varying levels of epistemic uncertainty by systematically varying the strength of the Bayesian priors. Moreover, by defining appropriate hyperpriors, Bayesian posterior inference can fully incorporate our epistemic uncertainty into the posterior distribution. Due to limited space, we highlight only the key pieces of the Bayesian framework here and leave the full specification to the Appendix.
5.1 Probability model
We model the assigned score $\hat{S}_{ik}^{(j)}$ by judge $j$ to candidate $k$’s answer to the $i$-th question, given its true score $S_{ik}^{*}$, as independent draws from a multinomial distribution with parameter $\theta_{S_{ik}^{*},k}^{(j)}$. [^2] Note that the model deliberately ignores the correlation between judge scores to avoid potential misspecification of complex correlation structures; we prove in the Appendix that the posterior inference procedure is still consistent. Marginalizing out the true latent scores, the likelihood of the observed data is
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big],$
where $n$ is the number of questions. To remove the influence of self-preference, note that we filter for $k\neq j$.
5.2 Prior for relaxing constancy
To relax the constancy assumptions (Assumptions 4 and 5), we model the prevalence of each judge-assigned score for a given candidate using random effects relative to a “base” score prevalence $\vec{\pi}_{k}$:
$\vec{\pi}_{k}^{(j)}=(1-W_{k}R_{j})\vec{\pi}_{k}+W_{k}R_{j}Z_{k}$
where $Z_{k}\sim\text{Dirichlet}(\delta)$ determines the direction of the deviation, $R_{j}\sim\text{Beta}(\omega J,J)$ controls the judge-specific magnitude, and $W_{k}\sim\text{Beta}(\omega K,K)$ controls the candidate-specific magnitude. The hyperparameter $\omega\in[0,\infty)$ controls the degree of constancy violation through the random effect magnitude, where $\omega=0$ enforces perfect constancy and increasing $\omega$ allows progressively larger violations. For instance, $\omega=1$ expects random effects to contribute half of the effect, while $\omega=8$ expects a contribution of 90%.
5.3 Prior for varying judge quality
Figure 5: Weight propagation framework for encoding judge quality priors in 3-level scoring, where outgoing edge weights must sum to 1. The probability of the judge assigning score $m_{2}$ when the true score is $m_{1}$, $\theta_{m_{1},m_{2}}$, equals the weighted sum of all its parent nodes per the incoming edge weights.
To construct a Bayesian prior over slopes between judge vertices, we parameterize the judge vertices in terms of a weight propagation graph (Figure B.10), where each element in the confusion matrix is an average of its parent nodes weighted by edge weights $\alpha$, i.e., $\theta_{\texttt{v}}^{(j)}=\sum_{\texttt{u}\rightarrow\texttt{v}}\theta_{\texttt{u}}^{(j)}\alpha_{\texttt{u}\rightarrow\texttt{v}}$. As such, the slopes between vertices are directly controlled in terms of edge weights. For each node v, we then place a Dirichlet prior over the outgoing edge weights $\alpha_{\texttt{v}}$ indexed by a single hyperparameter $\beta_{\max}$, where increasing $\beta_{\max}$ corresponds to more discriminative judges. For instance, in 3-level scoring, the Dirichlet prior for outgoing edges from node $\theta_{m_{1},m_{2}}$ has parameters $\vec{\beta}_{m_{1},m_{2}}=[1,1+\rho\cdot\beta_{\max},1]$ with $\rho\sim\text{Beta}(1,1)$. Full details are in the Appendix.
5.4 Conducting sensitivity analyses
Using this framework, we can develop a practical framework for conducting sensitivity analyses, to understand how epistemic uncertainty influences our confidence in the estimated rankings. In the absence of strong prior knowledge about LLM judge behavior, we recommend the following:
  1. 1.
    Start with conducting Bayesian inference with hyperparameters $\omega=0$ (perfect constancy) and $\beta_{\max}=5$ (moderate judge quality).
  2. 2.
    Estimate how rankings shift as the constancy assumption is relaxed, by increasing $\omega$ from 0 to 8 while holding $\beta_{\max}$ fixed.
  3. 3.
    Estimate how rankings shift as our prior over judge quality varies, by varying $\beta_{\max}$ from 0 to 20 while holding $\omega=0$.
If rankings remain stable across parameter variations, the conclusions are reliable despite epistemic uncertainty. Otherwise, one may need to inject additional prior knowledge (e.g., narrow the range of plausible hyperparameter values) or obtain gold-standard labels.
6 Experiments
Figure 6: Top: Sensitivity of estimated rankings when varying random effects (RE, left) and judge quality (right) hyperparameters, by plotting the correlation between the estimated ranking for each hyperparameter and their base values ($\omega=0,\beta_{\max}=0)$. Bottom: Candidates visualized on the probability simplex with judge configurations sampled from posterior (blue triangles), for MTBench (left) and Omni-MATH (right). We collapse 5-level ratings to 3 levels for visualization (see Appendix for mapping details).
Here we present experiments across various LLM benchmarks to assess the impact of epistemic uncertainty on ranking. First, using the geometric Bayesian framework, we conduct sensitivity analyses to determine the robustness of learned rankings. Second, we compare the overall performance of the Bayesian framework to existing methods in terms of coverage and correlation with ground truth rankings. We present our main experimental results below, with ablation studies, implementation details, and extended analyses in the Appendix.
We evaluate the Bayesian framework across three categories of benchmark datasets: (i) Verifiable tasks with 2-level scores: GPQA (Rein2023-dv) and MMLU Pro (Wang2024-ic) contain multiple-choice questions graded as correct/incorrect, with judges having an abstention option. (ii) Multi-level human-judged tasks: MTBench (Zheng2023-wr) evaluates multi-turn conversations while SummEval (Volske2017-bv) assesses summarizations. Both datasets are assessed on multi-level Likert scales, by LLMs as well as human experts. (iii) “Semi-verifiable” tasks: Omni-MATH (Gao2024-ec) contains mathematical reasoning problems with reference solutions, though there is no single ground-truth answer. LLM judges are asked to evaluate answers on a 3-level scale (correct, partial credit, incorrect).
We use a two-stage protocol to compare ranking methods. First, LLM judges evaluate candidates without access to ground truth, mirroring real-world usage. Second, we generate ground-truth scores for each answer by comparing against the correct multiple-choice answer on verifiable tasks and obtaining human-assigned scores for human-judged tasks. For semi-verifiable tasks, we rescore each candidate’s answer by providing the LLM judge with the provided reference answer. For GPQA, MMLU Pro, and Omni-MATH, we evaluated 19 contemporary models including Claude, GPT, Gemini, Llama, Mistral, and Qwen variants, with Claude 3.5 Haiku and GPT-4o mini serving as judges. These judges were deliberately selected to assess the performance of ranking methods when the judges are imperfect. For MTBench and SummEval, we utilize their existing candidates and judges.
6.1 Sensitivity analyses
Figure 7: Estimated rankings for the top 10 candidates on GPQA, presented in order by their true rankings. Unlike existing methods, Bayesian method fully incorporates epistemic uncertainty.
Table 1: Ranking method performance on benchmarks in terms of correlation and coverage of ground-truth ranks.
Method GPQA MMLU Omni-MATH SummEval MTBench Corr Cov Corr Cov Corr Cov Corr Cov Corr Cov Bayesian 0.881 (0.79, 0.90) 0.778 (0.68, 0.93) 0.862 (0.79, 0.88) 0.947 (0.95, 1.00) 0.696 (0.56, 0.75) 0.526 (0.38, 0.74) 0.888 (0.76, 0.91) 0.917 (0.67, 0.95) 1.000 (0.94, 1.00) 1.000 (1.00, 1.00) Bootstrap 0.847 (0.77, 0.89) 0.500 (0.56, 0.78) 0.826 (0.78, 0.87) 0.421 (0.47, 0.68) 0.682 (0.64, 0.69) 0.316 (0.26, 0.42) 0.885 (0.78, 0.90) 0.729 (0.65, 0.96) 0.943 (0.94, 1.00) 1.000 (1.00, 1.00) Bradley-Terry 0.847 (0.80, 0.88) 0.444 (0.40, 0.65) 0.842 (0.81, 0.87) 0.316 (0.33, 0.57) 0.682 (0.63, 0.69) 0.263 (0.21, 0.32) 0.902 (0.82, 0.91) 0.771 (0.61, 0.91) 0.943 (0.94, 1.00) 1.000 (0.67, 1.00) Simple Average 0.847 (0.77, 0.89) 0.167 (0.01, 0.28) 0.826 (0.78, 0.87) 0.105 (0.11, 0.21) 0.682 (0.64, 0.69) 0.158 (0.11, 0.21) 0.885 (0.78, 0.90) 0.271 (0.09, 0.54) 0.943 (0.94, 1.00) 0.667 (0.67, 1.00) Single Judge 0.817 (0.72, 0.87) 0.167 (0.01, 0.32) 0.825 (0.78, 0.87) 0.184 (0.11, 0.24) 0.632 (0.61, 0.67) 0.158 (0.11, 0.24) 0.787 (0.58, 0.86) 0.285 (0.06, 0.45) 0.971 (0.94, 1.00) 0.833 (0.67, 1.00)
For each benchmark dataset, we conducted Bayesian inference with random effects hyperparameter $\omega$ values of 0, 1, 2, 4, and 8, corresponding to expected random effects magnitudes of 0, 0.5, 0.7, 0.8 and 0.9 respectively. Figure 6 (left) shows that rankings remain remarkably stable across these variations for most datasets. The correlation between rankings at different $\omega$ values and the baseline rankings exceeds 0.95 for GPQA, MTBench, and SummEval, indicating that significant violations of the constancy assumption have minimal impact on the resulting rankings. Similarly, varying the judge quality hyperparameter $\beta_{\max}$ from 1 to 20 produces negligible effects on rankings across all datasets (Figure 6 right), with correlations consistently above 0.99. However, both MMLU Pro and Omni-MATH show greater sensitivity to the constancy assumption, with correlations between baseline rankings dropping to 0.77 and 0.86 respectively when $\omega$ is increased from 0 to 8. This sensitivity appears pronounced for MMLU Pro for $\omega>4$, highlighting uncertainty in its rankings because of potential variability in judge performance across candidates.
To understand why estimated rankings are more or less robust, we can visualize the candidates in the probability simplex against possible judge configurations sampled from the posterior distribution of the Bayesian method. Here we compare two contrasting datasets to illustrate why some benchmarks are robust to epistemic uncertainty while others are not. As shown in Figure 6 left, estimated rankings for MTBench are robust even when the constancy assumption is relaxed because (i) candidates are well-separated across the simplex, with clear performance tiers and (ii) the posterior judge configurations are relatively consistent with similar triangular shapes with minimal variability. In contrast, Figure 6 right shows that estimated rankings for Omni-MATH are more sensitive to epistemic uncertainty because (i) candidates cluster tightly in the central region with overlapping score distributions, making their relative ordering sensitive to judge vertex positions, and (ii) the posterior samples demonstrate considerable uncertainty in the judge configurations.
6.2 Comparative performance evaluation
We now evaluate the overall performance of the Bayesian framework against established baselines. To summarize the performance of the Bayesian procedure in a single metric, we fully incorporate epistemic uncertainty during posterior inference by defining a hyperprior over the hyperparameters varying the degree of constancy violation and judge quality. We compare against single judge scoring, score averaging across multiple judges (i.e., LLM judges as a jury), bootstrap confidence intervals (Xie2009-wv), and Bradley-Terry pairwise comparison models (Bradley1952-kj). Performance was evaluated in terms of Spearman correlation with the ground truth rankings and coverage rates of 95% credible/confidence intervals, averaged across candidates.
The Bayesian framework consistently achieves higher correlations and better coverage across all benchmark datasets (Table 1), with coverage rates improving by over 20 percentage points in some cases. This dramatic improvement in coverage reflects our theoretical finding that epistemic uncertainty is essential for proper uncertainty quantification. The correlation improvements, while more modest, demonstrate that incorporating judge uncertainty also benefits ranking estimates. Figure 7 illustrates how this coverage improvement is achieved: as seen on the GPQA dataset, the intervals from Bayesian inference are indeed significantly wider than those produced by comparator methods. Posterior distributions for the other datasets are given in the Appendix.
7 Discussion
This work leverages a novel geometric framework for understanding when gold-standard rankings can and cannot be recovered when one has access to imperfect LLM judges but no gold-standard labels. By representing judges and candidates on a probability simplex, we establish theoretical conditions for ranking identifiability and use these results to develop a sensitivity analysis approach. Our key insight is that epistemic uncertainty—uncertainty about judge quality—critically affects ranking reliability and is not considered by existing methods. The geometric perspective reveals that identifiability depends on scoring levels, judge consistency, and candidate separation in the simplex.
Appendix A Geometric Perspective: Additional details
A.1 Expected scores map to height in augmented space
A key geometric insight used throughout our analysis is that a candidate’s expected score corresponds to its height when the probability simplex is embedded in a higher-dimensional space. This result connects the abstract barycentric coordinates to the concrete ranking task.
Theorem 3 (Height Correspondence in Augmented Simplex).
Consider embedding the (M1)(M-1)-dimensional probability simplex into $M$-dimensional space by adding a “score” axis. Lift each judge vertex $\theta_{m}^{(j)}$ to position $(\theta_{m}^{(j)},m)$, placing it at height $m$. For any candidate $k$ with barycentric coordinates $\vec{\pi}_{k}=(\pi_{k,1},\ldots,\pi_{k,M})$, when positioned in the augmented space using these same coordinates, its height equals its expected score $\mathbb{E}[S_{k}^{*}]$.
Proof.
The candidate’s position in the original (M1)(M-1)-dimensional simplex is determined by its barycentric coordinates:
$\gamma_{k}^{(j)}=\sum_{m=1}^{M}\pi_{k,m}\theta_{m}^{(j)}$ (2)
where $\pi_{k,m}$ represents the weight given to judge vertex $\theta_{m}^{(j)}$.
To augment the space, we embed the simplex into $M$-dimensional space by adding a “score” coordinate. Each judge vertex $\theta_{m}^{(j)}$ is lifted to position $(\theta_{m}^{(j)},m)$, placing it at height equal to its score value $m$.
Crucially, when we position the candidate in this augmented space using the same barycentric coordinates $\pi_{k,m}$, we get:
$\text{Augmented position}=\sum_{m=1}^{M}\pi_{k,m}\left(\theta_{m}^{(j)},m\right)$ (3)
This convex combination yields:
  • Simplex coordinates: $\gamma_{k}^{(j)}=\sum_{m=1}^{M}\pi_{k,m}\theta_{m}^{(j)}$ (unchanged)
  • Height coordinate: $h_{k}=\sum_{m=1}^{M}\pi_{k,m}\cdot m$
The height $h_{k}=\sum_{m=1}^{M}\pi_{k,m}\cdot m$ is precisely the definition of the expected score $\mathbb{E}[S_{k}^{*}]$, where $\pi_{k,m}=\Pr(S_{k}^{*}=m)$.
Thus, the candidate’s height in the augmented space directly encodes its expected score, providing a visual representation where ranking candidates corresponds to comparing their heights. ∎
This geometric correspondence provides intuitive visualization: candidates automatically acquire heights equal to their expected scores when the simplex is embedded in augmented space using the barycentric coordinates. Ranking becomes a simple matter of comparing heights. Moreover, this representation naturally captures how uncertainty in judge vertex positions (epistemic uncertainty) translates to uncertainty in candidate heights and thus rankings.
A.2 Recoverability is dataset-specific
Figure A.8: Example where candidate rankings are likely identifiable even when there are 3+ levels, due to characteristics of the dataset. Candidates have assigned scores that are very different (mostly 1’s versus mostly 3’s).
There exist special cases where rankings become robust to judge quality assumptions despite the general non-identifiability result for 3+-level scoring systems. When candidates exhibit substantially different score distributions—manifesting as widely separated points on the probability simplex—their relative ordering becomes stable across a broad range of plausible judge configurations. See Figure A.8 as an example (Example 2 in Section 4.3).
Appendix B Bayesian model
B.1 Probability Model
For judge $j$ evaluating candidate $k$’s answer to the $i$-th question, we assume the assigned score $\hat{S}_{ik}^{(j)}$ given its true score $S_{ik}^{*}$ follows a multinomial distribution with parameter $\theta_{S_{ik}^{*},k}^{(j)}$ (Figure B.9). To maintain computational tractability while preserving the geometric structure identified in Section 3, we assume conditional independence across judges and questions. After marginalizing over the true latent scores, the likelihood of the observed data becomes:
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big],$ (4)
where $n$ is the number of questions and we exclude self-evaluations ($k\neq j$) to avoid self-preference bias.
Figure B.9: Plate diagram of the Bayesian model
Although judge scores are likely correlated in practice and (4) simplifies the true data generating mechanism, modeling the full joint distribution would increase complexity and require additional constancy assumptions without changing the identifiability results in Section 3. As we show in Theorem 5, for binary scoring systems, posterior inference under this model yields consistent estimators for prevalences and rankings even when correlation structure is ignored.
B.2 Random Effects Parameterization
B.2.1 Complete Hierarchical Specification
To relax the constancy assumption while maintaining computational tractability, we introduce random effects that allow judge-specific and candidate-specific deviations from base prevalences. The complete hierarchical model is:
$Z_{k}$ $\sim\text{Dirichlet}(\delta)\quad\text{(candidate-specific random direction)}$ (5)
$R_{j}$ $\sim\text{Beta}(\omega J,J)\quad\text{(judge-specific random effect magnitude)}$ (6)
$W_{k}$ $\sim\text{Beta}(\omega K,K)\quad\text{(candidate-specific random effect magnitude)}$ (7)
$\vec{\pi}_{k}^{(j)}$ $=(1-W_{k}R_{j})\vec{\pi}_{k}+W_{k}R_{j}Z_{k}\quad\text{(perturbed prevalences)}$ (8)
The perturbed prevalences $\vec{\pi}_{k}^{(j)}$ replace $\vec{\pi}_{k}$ in the likelihood (4). This parameterization ensures that $\vec{\pi}_{k}^{(j)}$ remains on the probability simplex for all valid parameter values.
B.2.2 Choice of Prevalence vs. Confusion Matrix Perturbation
We considered two approaches for introducing random effects: we can either allow for perturbations in judge performance (i.e. $\theta_{m,k}^{(j)}$ as perturbations of $\theta_{m}^{(j)}$) or score prevalences (i.e. $\vec{\pi}_{k}^{(j)}$ as perturbations of $\vec{\pi}_{k}$).
We chose prevalence perturbation for several reasons. For $M$-level scoring with $J$ judges and $K$ candidates, confusion matrix perturbation requires $O(JKM^{2})$ parameters while prevalence perturbation requires only $O(JKM)$. Fewer parameters reduce the risk of overfitting and model misspecification and enables faster convergence in MCMC sampling. Finally, deviations in prevalences directly correspond to changes in perceived candidate quality by different judges.
B.2.3 Setting $\delta$ for Detecting Specific Biases
The Dirichlet parameter $\delta$ controls the distribution of random directions and can be tuned to detect specific types of bias:
  • Uniform exploration: $\delta=[1,1,...,1]$ explores all directions equally
  • Score inflation detection: For 3-level scoring, $\delta=[1,4,10]$ places more weight on higher scores, detecting judges who systematically overrate candidates
  • Score deflation detection: For 3-level scoring, $\delta=[10,4,1]$ places more weight on lower scores, detecting overly critical judges
  • Central tendency bias: $\delta=[1,10,1]$ for 3-level scoring detects judges who avoid extreme scores
In our experiments, we use score inflation detection ($\delta=[1,4,10]$) as self-preference bias typically manifests as inflated scores for one’s own model family.
B.3 Judge Quality Prior Specification
The judge quality prior is implemented through a transition weight framework that parameterizes the confusion matrix columns (judge vertices) as weighted combinations of base positions. This section provides a detailed mathematical specification of the framework briefly introduced in Section 5.3 of the main paper. This approach allows us to encode beliefs about relative vertex positions while maintaining the probabilistic constraints of the confusion matrix.
B.3.1 Transition Weight Framework
We can parameterize the judge’s confusion matrix columns (vertices) through a directed acyclic graph where nodes represent confusion matrix entries and edges control how probability mass flows between them. This allows us to encode prior beliefs about judge quality through the edge weights while ensuring mathematical constraints are satisfied.
$\left\{\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\right\}_{(m_{1}^{\prime},m_{2}^{\prime})\text{ s.t. }(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\sim\mathrm{Dirichlet}(\vec{\beta}_{(m_{1},m_{2})})$
$\theta_{m_{1}^{\prime},m_{2}^{\prime}}^{(j)}=\sum_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\theta_{(m_{1},m_{2})}^{(j)}\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}$
Figure B.10: Transition weight framework for encoding judge quality priors in 3-level scoring. Nodes represent pairs $(m_{1},m_{2})$ where $m_{1}$ is the true score and $m_{2}$ is the assigned score. Edges show allowed transitions, with weights $\alpha$ drawn from Dirichlet priors parameterized by $\vec{\beta}_{(m_{1},m_{2})}$ (see Appendix B.3 for formal specification).
The formal specification of the transition weights is:
$\left\{\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\right\}_{(m_{1}^{\prime},m_{2}^{\prime})\text{ s.t. }(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\sim\text{Dirichlet}(\vec{\beta}_{(m_{1},m_{2})})$
$\theta_{m_{1}^{\prime},m_{2}^{\prime}}^{(j)}$ $=\sum_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\theta_{(m_{1},m_{2})}^{(j)}\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}$
In this notation, each node $(m_{1},m_{2})$ in the transition graph represents an entry in the confusion matrix: the probability of assigning score $m_{2}$ when the true score is $m_{1}$. The transition weights $\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}$ control how probability mass flows from parent nodes to child nodes. All outgoing weights from any parent node must sum to one, ensuring the result remains a valid probability distribution. The final confusion matrix entry $\theta_{m_{1}^{\prime},m_{2}^{\prime}}^{(j)}$ for judge $j$ is computed as a weighted average of its parent nodes’ values, weighted by the incoming edge weights.
For example, in Figure B.10, the probability $\theta_{2,2}$ (assigning score 2 when true score is 2) receives contributions from both $\theta_{1,1}$ and $\theta_{1,2}$ through their respective edge weights.
B.3.2 Parameterization for 3-Level Scoring
For a 3-level scoring system, the Dirichlet parameters $\vec{\beta}_{(m_{1},m_{2})}$ are specified as follows:
$\vec{\beta}_{1,1}^{(j)}$ $=[1,1+\rho^{(j)}\beta_{\max},1]$
$\vec{\beta}_{1,2}^{(j)}$ $=[{\beta}_{1,1,2}^{(j)},{\beta}_{1,1,3}^{(j)}]$
$\vec{\beta}_{2,1}^{(j)}$ $=[1,1,1+\rho^{(j)}\beta_{\max}]$
$\vec{\beta}_{2,2}^{(j)}$ $=[{\beta}_{2,1,2}^{(j)},{\beta}_{2,1,3}^{(j)}],$
where:
  • $\beta_{\max}$ is the global hyperparameter controlling prior strength about judge quality
  • $\rho^{(j)}\sim\text{Beta}(1,1)$ allows judge-specific variation in discrimination ability (drawn from a uniform hyperprior on $[0,1]$)
  • The structure ensures higher weight on correct identification (diagonal elements) as $\beta_{\max}$ increases
Note that $\alpha_{1,3}$ and $\alpha_{2,3}$ are deterministically set equal to one as these represent terminal nodes in the transition graph with no outgoing edges.
Intuitively, when $\beta_{\max}=0$, all Dirichlet parameters equal 1 (uniform prior), allowing maximum flexibility in judge behavior. As $\beta_{\max}$ increases, the parameters $1,1+\rho^{(j)}\beta_{\max},1]$ place more weight on the middle transition, encoding our belief that better judges more consistently assign higher scores when the true score increases.
B.3.3 Enforcing Monotonicity Through Graph Structure
The transition graph topology automatically enforces Assumption 3 (monotonicity) by restricting which edges exist. Specifically, edges from node $(m_{1},m_{2})$ to $(m_{1}^{\prime},m_{2}^{\prime})$ only exist when:
  • $m_{1}^{\prime}=m_{1}+1$ (moving to the next true score level)
  • $m_{2}^{\prime}\geq m_{2}$ (assigned score cannot decrease as true score increases)
This structural constraint ensures that judges cannot be “label-flippers” who systematically assign lower scores to better answers, without requiring explicit inequality constraints during MCMC sampling.
B.3.4 Extension to $M$-Level Scoring
For general $M$-level scoring systems, the framework extends naturally:
  • The graph has $M(M+1)/2$ nodes representing all valid $(m_{1},m_{2})$ pairs where $m_{2}\geq m_{1}$
  • Edge constraints follow the same monotonicity rules
  • Dirichlet parameters follow the pattern: higher values along paths that preserve or increase the assigned score relative to the true score
The specific parameterization for $M>3$ follows similar principles, with $\beta_{\max}$ controlling the concentration around correct identification. See the accompanying code for exact implementation details for different values of $M$.
Appendix C Implementation details
Bayesian inference: We conduct posterior inference using Hamiltonian Monte Carlo (HMC), coded up using Stan (Stan_Development_Team2021-id). HMC ran with 4 chains, each running 1000 iterations for warmup and another 1000 for sampling.
Abstention: In practice, the judge may not always be sure what score to assign. So we give the judge the option to abstain when there are only two levels, e.g. correct or not.
Appendix D Experiment configuration
D.1 Benchmarks
  • GPQA (Rein2023-dv): Questions resistant to simple internet searches across STEM domains, stratified by difficulty (undergraduate, graduate, post-graduate).
  • MMLU Pro (Wang2024-ic): Enhanced professional knowledge questions across 16 domain-specific subcategories from natural sciences, social sciences, and humanities.
  • MTBench (Zheng2023-wr): A conversational benchmark evaluating single- and multi-turn dialogue capabilities across diverse scenarios (creative writing, reasoning, coding, mathematics, role-playing). Human judges rated responses on a 10-point scale. For LLM judges, we ask for ratings on a simplified 5-point scale.
  • TLDR (aka SummEval) (Fabbri2021-ie): A summarization benchmark where models condense news articles into concise summaries, with human ratings across four dimensions (relevance, consistency, fluency, coherence) on 5-point scales.
  • Omni-MATH (Gao2024-ec): A benchmark focusing on high-difficulty competition-level problems from International and National Olympiads. These problems present particular challenges for automated evaluation as solutions vary significantly in approach, notation, and presentation; multiple valid solution paths may exist for the same problem; and partial correctness must be assessed along multiple dimensions.
D.2 LLM judges
We use two LLM Claude 3.5 Haiku (anthropic/claude-3-5-haiku-20241022) and GPT-4o Mini (gpt-4o-mini-2024-07-18) for all benchmark datasets but TLDR. For TLDR, we use the provided judge-assigned scores from older LLMs: GPT-4-0314, GPT-3.5-Turbo-0301, and Llama-2-70b-chat-hf.
To mitigate position bias in evaluation, we randomized the order of candidate responses presented to judges.
D.3 LLM candidates
For consistent comparison across datasets, we evaluated the set of candidate LLMs shown in Table 2.
Table 2: LLM candidates evaluated across all experimental settings
Model Family Model Name Version/Date Experiments Anthropic Claude 3.5 Haiku claude-3-5-haiku-20241022 GPQA, MMLU Pro, Omni-MATH Anthropic Claude 3.5 Sonnet claude-3-5-sonnet-20241022 GPQA, MMLU Pro, Omni-MATH Anthropic Claude 3.7 Sonnet claude-3-7-sonnet-20250219 GPQA, MMLU Pro, Omni-MATH DeepSeek DeepSeek V3 deepseek-v3 GPQA Google Gemini 1.5 Flash gemini-1.5-flash-002 GPQA, MMLU Pro, Omni-MATH Google Gemini 1.5 Pro gemini-1.5-pro-002 GPQA, MMLU Pro, Omni-MATH Google Gemini 2.0 Flash gemini-2.0-flash-001 GPQA, MMLU Pro, Omni-MATH Google Gemini 2.0 Flash Lite gemini-2.0-flash-lite-preview-02-05 GPQA, MMLU Pro, Omni-MATH Meta Llama 3.1 405B llama-3.1-405b-instruct-turbo GPQA, MMLU Pro, Omni-MATH Meta Llama 3.1 70B llama-3.1-70b-instruct-turbo GPQA, MMLU Pro, Omni-MATH Meta Llama 3.1 8B llama-3.1-8b-instruct-turbo GPQA, MMLU Pro, Omni-MATH Meta Llama 4 Maverick 17B llama-4-maverick-17b-128e-instruct-fp8 MMLU Pro, Omni-MATH Meta Llama 4 Scout 17B llama-4-scout-17b-16e-instruct MMLU Pro, Omni-MATH Mistral AI Mistral 7B mistral-7b-instruct-v0.3 GPQA Mistral AI Mixtral 8x22B mixtral-8x22b-instruct-v0.1 GPQA Mistral AI Mixtral 8x7B mixtral-8x7b-instruct-v0.1 GPQA OpenAI GPT-4.1 gpt-4.1-2025-04-14 MMLU Pro, Omni-MATH OpenAI GPT-4.1 mini gpt-4.1-mini-2025-04-14 MMLU Pro, Omni-MATH OpenAI GPT-4.1 nano gpt-4.1-nano-2025-04-14 MMLU Pro, Omni-MATH OpenAI GPT-4o gpt-4o-2024-11-20 GPQA, MMLU Pro, Omni-MATH OpenAI GPT-4o mini gpt-4o-mini-2024-07-18 GPQA, MMLU Pro, Omni-MATH Qwen Qwen 2.5 72B qwen2.5-72b-instruct-turbo GPQA, MMLU Pro, Omni-MATH Qwen Qwen 2.5 7B qwen2.5-7b-instruct-turbo GPQA, MMLU Pro, Omni-MATH
For the MTBench dataset (Zheng2023-wr), we assess the candidates in the provided dataset: GPT-4-0613, Claude-1, Llama-2-13B-Chat, Vicuna-13B, and Alpaca-13B. Similarly, for the TLDR (SummEval) benchmark (Fabbri2021-ie), we assess the 12 provided language models.
D.3.1 Comparison Methods
We compare our Bayesian adjudication framework against the following baseline methods:
Simple Averaging: This approach computes the mean score for each candidate across all evaluations and determines rankings based on these averages. It treats each judge’s assessment with equal weight and assumes judge scores accurately reflect true performance.
Single Judge Aggregation: This approach collapses distinctions between multiple judges, treating all evaluations as if they came from a single judge. It computes the mean score for each candidate across all judge evaluations, ignoring judge identity.
Simple Averaging with Bootstrap Confidence Intervals: This comparator uses the bootstrap approach for population ranks proposed in Xie2009-wv to generate confidence intervals for the simple averaging estimate.
Pairwise Comparison Approach: As a representative pairwise comparison method, we an extension of the Bradley-Terry model with ties (Rao1967-fv), which estimates candidate ability parameters based on win-loss-tie patterns in pairwise evaluations. We calculate confidence intervals for this approach using bootstrap.
Prediction-Powered Inference (PPI): For benchmarks with ground truth labels, we implement PPI (Angelopoulos2023-tv) as an alternative ranking method. PPI uses a small labeled dataset to calibrate predictions from LLM judges on a larger unlabeled dataset, providing statistically valid confidence intervals for candidate rankings. Our implementation: (i) randomly partitions questions into labeled and unlabeled sets (using 5% or 10% labeled fractions), (ii) applies the PPI mean estimator using the ppi_py library with judge scores converted to a 0-1 scale, and (iii) generates confidence intervals through bootstrap resampling (300 iterations) while maintaining the fixed labeled/unlabeled partition. We apply PPI to GPQA, MMLU Pro, and Omni-MATH benchmarks where ground truth answers are available. Results for different labeled data fractions are presented in Section E.1.
To assess the performance of our framework against these baselines, we use several evaluation metrics. We measure the accuracy of ranking estimates using Spearman’s rank correlation with ground truth rankings when available. The quality of uncertainty quantification is evaluated through coverage rates of 95% credible/confidence intervals. This combination of metrics allows us to evaluate both the point estimate accuracy and the calibration of uncertainty in our framework’s predictions, providing a holistic assessment of each methods performance in the evaluation tasks.
D.4 Prompts
Below are the prompts used for evaluating candidate answers in each of our experiments. Where possible, we use structured generation to produce JSON structured outputs.
D.4.1 Binary Verification Judge Prompt
For the GPQA and MMLU Pro dataset experiments where we have multiple-choice questions with ground truth answers, we use the following prompt:
You are evaluating candidate answers to a multiple-choice question.
- Consistency: How well the candidates explanation aligns with their final multiple choice selection (1-5 scale).
* 1 = The explanation contradicts the selected answer
* 2 = Major disconnects between explanation and selected answer
* 3 = Explanation partially supports the answer with some inconsistencies
* 4 = Explanation mostly supports the answer with minor inconsistencies
* 5 = Explanation perfectly aligns with and justifies the selected answer
- Accuracy: Did the candidate select the correct answer choice? (-1 = no, 1 = yes, 0 = unsure)
* Provide a concise explanation referencing key facts or reasoning that makes the answer correct or incorrect
<QUESTION>
[[question]]
</QUESTION>
[[candidates_section]]
Respond with a JSON object containing evaluations for all candidates and ensure that your JSON response:
- Uses the exact structure provided below
- Includes only the evaluation without additional preamble or commentary
- Properly escapes any special characters in the reasoning strings
"evaluations" : [
"model_id": "1",
"consistency": {
"reasoning": str,
"score": int
"accuracy": {
"reasoning": str,
"score": int
In the actual implementation, the [[candidates_section]] placeholder is dynamically populated with candidate answers and explanations using the following format:
<CANDIDATE #{i} ANSWER>
{candidate.get(answer_key)}
</CANDIDATE ANSWER>
<CANDIDATE #{i} EXPLANATION>
{rationale}
</CANDIDATE #{i} EXPLANATION>
where {i} is the candidate number, {candidate.get(answer_key)} retrieves the candidate’s multiple-choice answer, and {rationale} contains their explanation for that answer.
D.4.2 Human Judgement Prompts
MTBench Judge Prompts:
For MTBench, we use a prompt similar to the original prompt given to human judges as described in (Zheng2023-wr) for the single-turn and two-turn cases. Similar to the binary tasks, we replace the [[question]] and [[candidates_section]] placeholders with the prompt and the conversation history using the following format:
Please act as an impartial judge and evaluate the quality of the responses provided by AI assistants to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants.Be as objective as possible. Rate the response on a scale of 1 to 5 (1=Very Bad, 5=Very Good), along with the reasoning.
Initial User Question:
<PROMPT>
[[question]]
</PROMPT>
Assistant Conversation(s):
[[candidates_section]]
"evaluations" : [
"model_id": "1",
"overall": {
"reasoning": str,
"score": int
"model_id": "2",
"overall": {
"reasoning": str,
"score": int
"model_id": "3",
"overall": {
"reasoning": str,
"score": int
The candidate presentation format differs between single-turn and multi-turn evaluations. For single-turn interactions, we present only the initial response:
<CANDIDATE #{i}>
{response1}
</CANDIDATE #{i}>
For two-turn interactions, we present the complete conversation history with delineation between turns:
<CANDIDATE #{i}>
<TURN 1>
[User Prompt]
{prompt1}
[Assistant Response]
{response1}
</TURN 1>
<TURN 2>
[User Prompt]
{prompt2}
[Assistant Response]
{response2}
</TURN 2>
</CANDIDATE #{i}>
In cases where a candidate fails to respond to the second turn, we explicitly note this absence:
<TURN 2>
[No second turn response provided]
</TURN 2>
TLDR (SummEval) Judge Prompts:
For the TLDR (SummEval) dataset, we use the same evaluation framework as in (Fabbri2021-ie) that assesses news article summarization quality across four distinct criteria: relevance, consistency, fluency, and coherence. Each dimension is evaluated on a 5-point Likert scale with specific definitions provided to ensure consistent interpretation:
Instructions: In this task you will evaluate the quality of summaries written for a news article. You will be shown the original article and [[num_candidates]] candidate summaries.
To correctly solve this task, follow these steps:
1. Carefully read the original news article provided below.
2. Read the candidate summaries presented in the <CANDIDATE #i ANSWER> sections.
3. Rate each summary on a scale from 1 (very low) to 5 (very high) based on its relevance, consistency, fluency, and coherence. Note that summaries that are very similar on an axis may receive the same score.
Definitions:
* Relevance: The rating measures how well the summary captures the key points of the article. Summaries in which all and only the important aspects are contained will receive the highest rating.
* Consistency: The rating measures whether the facts in the summary are consistent with the facts in the original article. The summary should stay true to the facts reported and not make up untrue information.
* Fluency: This rating measures the quality of individual sentences: are they well-written and grammatically correct?
* Coherence: This rating measures the quality of all sentences collectively: do they fit together and sound natural? Consider the quality of the summary as a whole.
Original news article:
[[question]]
Candidate Summaries:
[[candidates_section]]
Now provide your scores in the following JSON format. Ensure your response is a single JSON object, starting with {{ and ending with }}, and includes evaluations for all [[num_candidates]] candidates:
"evaluations": [
// Evaluation for Candidate #1
"model_id": "1", // Corresponds to Candidate #1
"relevance": {{
"reasoning": "Provide your reasoning for the relevance score here.",
"score": int // Score from 1 to 5
"consistency": {{
"reasoning": "Provide your reasoning for the consistency score here.",
"score": int // Score from 1 to 5
"fluency": {{
"reasoning": "Provide your reasoning for the fluency score here.",
"score": int // Score from 1 to 5
"coherence": {{
"reasoning": "Provide your reasoning for the coherence score here.",
"score": int // Score from 1 to 5
// Add evaluations for Candidate #2, #3, up to #[[num_candidates]] following the same structure
// Example for Candidate #2:
"model_id": "2", // Corresponds to Candidate #2
"relevance": {{
"reasoning": "…",
"score": int
"consistency": {{
"reasoning": "…",
"score": int
"fluency": {{
"reasoning": "…",
"score": int
"coherence": {{
"reasoning": "…",
"score": int
// other candidates
Here we fill in the [[question]] and [[candidates_section]] placeholders with the news article and the candidate summaries with the same format as the other datasets.
D.4.3 Semi-verifiable task Judge Prompt
For the Omni-MATH dataset experiments, we implement a two-stage evaluation process that addresses the unique challenges of assessing mathematical reasoning when multiple solution paths may be valid. This approach allows us to measure both standalone solution quality and alignment with reference solutions.
Stage 1: Evaluation Without Ground Truth:
In the first stage, judge LLMs evaluate candidate solutions based solely on their mathematical correctness without access to reference answers. This mimics how human experts might evaluate mathematical work without preconceived notions of the "correct" approach. The prompt emphasizes accuracy assessment on a 3-point scale (-1 for incorrect, 0 for partially correct, 1 for correct), with explicit instructions to use the middle category sparingly. The full prompt is as follows:
Instructions: Evaluate the quality of candidate answers to mathematical questions. You will be shown the original question and [[num_candidates]] candidate answers.
To correctly solve this task, follow these steps:
1. Carefully read the original question to understand what is being asked.
2. Read each candidate answer carefully.
3. Rate each answer according to the criteria below based on general mathematical knowledge and reasoning.
4. Provide clear justification for each score with specific references to the candidates answer.
Rate each answer using the following criteria:
### Accuracy Assessment (1 for correct, 0 for partially correct/borderline, -1 for incorrect)
Based on your mathematical knowledge, how accurate is the candidate answer? Strive to categorize answers as either Correct (1) or Incorrect (-1). Reserve the Partially Correct/Borderline (0) score for answers that contain significant correct elements but also notable errors or omissions, making a definitive Correct/Incorrect judgment difficult, or for answers that are technically correct but incomplete in a way that affects the final conclusion.
* 1 (Correct): The answer is mathematically sound, reaches a valid conclusion, and is substantially free of errors.
* 0 (Partially Correct / Borderline): The answer contains significant correct elements but also notable errors or omissions preventing a clear "Correct" score OR the answer is technically correct but misses key steps or context, making it significantly less complete. Use this score sparingly.
* -1 (Incorrect): The answer contains significant mathematical errors or reaches an incorrect conclusion.
Question:
[[question]]
Candidates Summaries:
[[candidates_section]]
Respond with a JSON object containing evaluations for all candidates and ensure that your JSON response:
- Uses the exact structure provided below
- Includes only the evaluation without additional preamble or commentary
- Properly escapes any special characters in the reasoning strings
- Always output the reasoning before providing a final score
"evaluations" : [
"model_id": "1",
"accuracy": {
"reasoning": str,
"score": int
Stage 2: Evaluation With Ground Truth Reference:
In the second stage, judge LLMs re-evaluate candidate solutions with access to reference solutions. This provides a benchmark for assessing alignment with established approaches while still allowing for alternative valid solution paths. The prompt maintains the same 3-point scale but refocuses the evaluation criteria on comparison with the reference solution. The full prompt is as follows:
Instructions: Evaluate the quality of candidate answers to mathematical questions. You will be shown the original question, the ground truth reference answer, and [[num_candidates]] candidate answers.
To correctly solve this task, follow these steps:
1. Carefully read the original question.
2. Carefully read the ground truth reference answer to understand the correct approach and solution.
3. For each candidate answer:
- Read the entire response
- Evaluate it against the ground truth reference answer
- Score it according to the criteria below
- Provide clear justification for each score with specific references to both the candidate answer and ground truth
Rate each answer using the following criteria relative to the ground truth reference answer:
### Accuracy Assessment (1 for correct, 0 for partially correct/borderline, -1 for incorrect)
Based on the reference answer, how accurate is the candidate answer? Strive to categorize answers as either Correct (1) or Incorrect (-1). Reserve the Partially Correct/Borderline (0) score for answers that contain significant correct elements but also notable errors or omissions, making a definitive Correct/Incorrect judgment difficult, or for answers that are technically correct but incomplete in a way that affects the final conclusion compared to the reference.
* 1 (Correct): The answer reaches the same mathematical conclusion as the reference answer (even if using a different valid approach) and is substantially free of errors.
* 0 (Partially Correct / Borderline): The answer contains significant correct elements but also notable errors or omissions preventing a clear "Correct" score OR the answer is technically correct but misses key steps or context provided in the reference, making it significantly less complete. Use this score sparingly.
* -1 (Incorrect): The answer reaches a different conclusion from the reference answer or contains significant mathematical errors that invalidate the result.
Question:
[[question]]
Ground Truth Reference Answer:
[[ground_truth_answer]]
Candidates Summaries:
[[candidates_section]]
Respond with a JSON object containing evaluations for all candidates and ensure that your JSON response:
- Uses the exact structure provided below
- Includes only the evaluation without additional preamble or commentary
- Properly escapes any special characters in the reasoning strings
- Always output the reasoning before providing a final score
"evaluations" : [
"model_id": "1",
"accuracy": {
"reasoning": str,
"score": int
Combined Analysis:
This two-stage evaluation approach enables separate analyses of intrinsic solution quality and reference alignment. For our primary analysis, we use the second-stage (reference-based) evaluations as pseudo-ground truth when comparing our Bayesian ranking methods against baselines. The first-stage evaluations provide valuable insights into the judge’s standalone mathematical reasoning capabilities, allowing us to assess how often judges can identify correct solutions without reference answers. We hypothesize that the delta between stage one and stage two evaluations provides a useful signal about the difficulty of problems and the capabilities of both candidate models and judge models.
In the actual implementation, the [[question]], [[ground_truth_answer]], and [[candidates_section]] placeholders are dynamically populated with the mathematical problem statement, reference solution, and candidate solutions, respectively. The candidate solutions are presented in the same format as in the other experimental settings.
D.5 MTBench self-preference
Table 3 presents the frequency of scores (ranging from 1, low, to 5, high) assigned by the two LLM judges—Claude 3.5 Haiku and GPT-4o mini—to various candidate LLMs based on their responses to two-turn questions from the MTBench dataset. The results suggest potential self-preference biases in the LLM-based evaluations. The Claude 3.5 Haiku judge awarded its predecessor, Claude v1, a high frequency of top scores (48 instances of ’5’), surpassing other models like GPT-4 (40 instances of ’5’). The GPT-4o mini judge assigned an equal number of perfect ’5’ scores (50 instances each) to both its own family model, GPT-4, and to Claude v1. While GPT-4o mini gives the same number of top scores to GPT-4 and Claude, the scoores from the Claude judge suggests that models tend to receive more favorable evaluations from judges within the same model family, highlighting a self-preference.
Table 3: Distribution of scores (1-5) assigned by Claude 3.5 Haiku and GPT-4o mini judges to candidate LLMs on the MTBench two-turn benchmark. Cell values represent the frequency of each score.
Claude 3.5 Haiku GPT-4o mini
Model Name 1 2 3 4 5 1 2 3 4 5
Llama 13B 29 32 11 5 0 43 24 15 0 0
Alpaca 13B 3 33 33 8 3 17 25 29 10 0
Vicuna 13B v1.2 1 7 24 27 19 7 12 19 25 18
Claude v1 0 3 2 21 48 1 6 6 18 50
GPT-3.5 Turbo 0 2 11 38 20 1 7 14 23 37
GPT-4 0 1 6 26 40 0 4 8 20 50
D.6 Score Mapping for Simplex Visualizations
For the probability simplex visualizations of 3+-level Likert scale datasets, we map the original 5-point scales to 3-point scales to enable more interpretable representations within the simplex. The mapping functions $f:\{0,1,2,3,4,5\}\mapsto\{1,2,3\}$ are defined as follows:
TLDR mapping:
$f_{\text{TLDR}}(s)=\begin{cases}1,&\text{if }s\in\{1,2\}\\ 2,&\text{if }s\in\{0,3,4\}\\ 3,&\text{if }s=5\end{cases}$
MTBench mapping:
[…]
The major difference between these mapping is the treatment of score 4. For TLDR, we group it with score 3 in the middle category, while for MTBench, we group it with score 5 in the top category. These different groupings reflect the empirical distribution patterns observed in each dataset. MTBench evaluations exhibit more separation between high-performing candidates, while TLDR shows finer distinctions between middle performing candidates.
Appendix E Additional Results
E.1 Prediction-Powered Inference Results
Table 4 presents ranking results using Prediction-Powered Inference (PPI) across GPQA, MMLU Pro, and Omni-MATH benchmarks with varying amounts of labeled data (5% and 10%). PPI demonstrates strong performance, achieving high correlations with true rankings and consistently strong coverage rates. These results confirm that PPI is highly effective at quantifying uncertainty when calibrated with even small amounts of labeled data, providing reliable confidence intervals while maintaining competitive ranking accuracy. The method’s strength lies in its ability to leverage limited ground truth to correct for judge bias, making it an excellent choice when some labeled examples are available. In contrast, the Bayesian framework provides robust uncertainty quantification when gold standard labels are unavailable, addressing the complementary scenario where obtaining any ground truth is impractical or infeasible.
Table 4: Performance of PPI ranking method across benchmark datasets with different labeled fractions. Results show Spearman correlation (Corr) and coverage rates (Cov) with 95% confidence intervals.
Labeled Fraction GPQA MMLU Pro Omni-MATH Corr Cov Corr Cov Corr Cov 5% 0.773 (0.445, 0.889) 0.889 (0.778, 1.000) 0.902 (0.691, 0.919) 0.947 (0.749, 1.000) 0.789 (0.486, 0.905) 0.895 (0.789, 1.000) 10% 0.875 (0.529, 0.923) 0.944 (0.790, 1.000) 0.922 (0.782, 0.929) 1.000 (0.737, 0.947) 0.737 (0.621, 0.929) 0.895 (0.801, 1.000)
E.2 Posterior Distribution Plots
Figures E.11-E.13 present posterior distributions for candidate rankings across additional benchmark datasets, complementing the GPQA results shown in the main text. These visualizations reveal consistent patterns in how different methods quantify uncertainty.
MMLU Pro and Omni-MATH (FigureE.11): Both datasets exhibit similar characteristics to GPQA, with the Bayesian method producing notably wider credible intervals that capture epistemic uncertainty about judge quality. The increased width is particularly pronounced for middle-ranked candidates, where judge disagreement tends to be highest. For Omni-MATH specifically, we observe greater overall uncertainty across all methods, consistent with the sensitivity analysis showing this dataset’s vulnerability to constancy violations.
Figure E.11: Posterior distributions for candidate rankings on MMLU Pro (left) and Omni-MATH (right), with candidates ordered by true ranking (best at top).
SummEval (Figure E.13): Across the four evaluation dimensions (coherence, consistency, fluency, and relevance), the Bayesian approach maintains wider intervals while preserving accurate point estimates. The coherence and consistency dimensions show tighter clustering of posterior distributions compared to fluency and relevance, suggesting these aspects may be more reliably evaluated by LLM judges. The top-ranked candidates (M22, M23, M17) show relatively narrow intervals across all methods, indicating strong agreement on the best performers.
Figure E.12: Posterior distributions for candidate rankings on MTBench, with candidates ordered by true ranking (best at top).
MTBench (Figure E.12): The MTBench dataset exhibits tight posterior distributions across all methods, indicating strong judge consensus. Even the Bayesian approach produces relatively narrow credible intervals, suggesting low epistemic uncertainty for this dataset. This aligns with our sensitivity analysis showing MTBench maintains stable rankings under relaxed constancy assumptions, representing a case where judge-based evaluation proves particularly reliable.
A key pattern across all datasets is that while the three methods generally agree on median rankings, they diverge substantially in uncertainty quantification. The Bootstrap and Bradley-Terry methods produce similar narrow confidence intervals that account only for sampling variability. In contrast, the Bayesian framework’s wider intervals reflect both aleatoric and epistemic uncertainty, explaining its superior coverage rates reported in main text. This difference is most pronounced for candidates with intermediate performance levels, where judge quality uncertainty has the greatest impact on ranking uncertainty.
Figure E.13: Posterior distributions for candidate rankings on TLDR/SummEval across four evaluation dimensions: coherence (top left), consistency (top right), fluency (bottom left), and relevance (bottom right). Candidates are ordered by true ranking (best at top).
E.3 Robustness to Self-Preference Bias
While our main experiments demonstrate the framework’s ability to handle naturally occurring judge biases (e.g., MTBench self-preference in Table 3), we conducted additional controlled experiments to systematically evaluate robustness to self-preference bias.
Experimental Design: Using the GPQA dataset, we artificially induced self-preference bias by manipulating judge-assigned scores when judges evaluated candidates from the same model family. Specifically, when Claude judges evaluated Claude candidates or GPT judges evaluated GPT candidates, we upward-shifted the assigned score with probability 0.8. This creates a strong but realistic bias pattern similar to that observed in real evaluations.
Bias Injection Protocol:
  • For binary scores (correct/incorrect), incorrect answers were changed to correct with probability 0.8
  • For abstentions, we converted them to correct with probability 0.8
  • Non-self evaluations remained unchanged
This manipulation creates a challenging scenario where traditional averaging methods would systematically overestimate the performance of judge-candidates.
Table 5: Performance under simulated self-preference bias on GPQA dataset
Method Correlation Coverage
Bayesian with self-preference adjustment 0.811 0.852
Bayesian without adjustment 0.701 0.611
Bootstrap 0.680 0.519
Bradley-Terry 0.673 0.482
Simple Average 0.680 0.056
Single Judge 0.683 0.148
The results demonstrate that the Bayesian framework with self-preference adjustment effectively corrects for the induced bias, achieving correlation and coverage rates comparable to the unbiased setting shown in the main text. Without adjustment, all methods suffer substantial performance degradation, with the Bayesian method’s correlation dropping from 0.881 to 0.701.
Appendix F Proofs
We present the proofs for the main results in the paper. For the identifiability results, recall that the assumptions are as follows:
Assumption 4.
Judge $\hat{s}_{j}$ satisfies “strong constancy” if its confusion matrix is the same across all $K$ candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for $k=1,\cdots,K$.
Assumption 5.
Judge $\hat{s}_{j}$ satisfies “moderate constancy” if its confusion matrix is the same for all non-self candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for all $k\neq j$.
Assumption 6.
The $j$-th judge’s probability of assigning the lowest score when the true score is equal to $m$ decreases with respect to $m$.
F.1 Proof for Theorem 1
Proof for Theorem 1(i).
By Assumption 4, we have
${{\theta_{1}}}\;=\;\Pr\bigl(\hat{S}_{k}=1\mid S_{k}^{*}=1\bigr),\qquad{{\theta_{0}}}\;=\;\Pr\bigl(\hat{S}_{k}=1\mid S_{k}^{*}=0\bigr)$
for all $k=1,\cdots,K$. Moreover, by Assumption 6, we have that $\theta_{1}>\theta_{0}$.
For candidate $k$, let
$\pi_{k}\;=\;\Pr\bigl(S_{k}^{*}=1\bigr),\qquad\gamma_{k}\;=\;\Pr\bigl(\hat{S}_{k}=1\bigr).$
Then
$\gamma_{k}=\;\pi_{k}\,{{\theta_{1}}}\;+\;(1-\pi_{k})\,{{\theta_{0}}}.$
For any set of $K$ candidates with non-equal prevalences of the judge-assigned scores, suppose we have that WLOG that the ordering indices $k_{1},\cdots,k_{K}$ are such that $\gamma_{k_{1}}>\gamma_{k_{2}}>\cdots>\gamma_{k_{K}}$. This implies that
$\bigl[{{\theta_{0}}}+\pi_{k_{1}}({{\theta_{1}}}-{{\theta_{0}}})\bigr]\;>\;\bigl[{{\theta_{0}}}+\pi_{k_{2}}({{\theta_{1}}}-{{\theta_{0}}})\bigr]\;>\;\cdots\;>\;\bigl[{{\theta_{0}}}+\pi_{k_{K}}({{\theta_{1}}}-{{\theta_{0}}})\bigr]$
which implies that
$\pi_{k_{1}}>\pi_{k_{2}}>\cdots>\pi_{k_{K}}.$
Proof for Theorem 1(ii).
WLOG, let the candidates who are not also judges (referred to as core candidates) have indices $k=3,4,\cdots,J$. Let the two LLMs who are both judges and candidates be $k=j\in\{1,2\}$ (reffered to as judge-candidates).
Per Assumption 5, we have for each judge $j=1,2$ that there exists $\theta_{1}^{(j)}>\theta_{0}^{(j)}$ such that
${{\theta_{1}^{(j)}}}\;=\;\Pr\bigl(\hat{S}_{k}=1\mid S_{k}^{*}=1\bigr),\qquad{{\theta_{0}^{(j)}}}\;=\;\Pr\bigl(\hat{S}_{k}=1\mid S_{k}^{*}=0\bigr),$
for all core candidates $k$.
Applying our result for Theorem 1(i), we can rank core candidates $k=3$ and $k=4$ using their ranking on the line segment from $\theta_{1}^{(1)}$ to $\theta_{0}^{(1)}$ (or the line segment from $\theta_{1}^{(2)}$ to $\theta_{0}^{(2)}$). Then all core-candidates $k$ can then be jointly ranked (including $k=3,4$) by assessing:
$\frac{\gamma^{(1)}_{k}-\gamma^{(1)}_{k}}{\gamma^{(1)}_{4}-\gamma^{(1)}_{3}}=\frac{\pi_{k}-\pi_{3}}{\pi_{4}-\pi_{3}}.$
We can rank judge-candidates in a similar way, using their non-self-judged position. Specifically, for candidate $k=1$, we use its score distribution from judge $j=2$, i.e. $\gamma^{(2)}_{1}=\Pr(\hat{S}_{1}^{(2)}=1)$, relative to candidates $k=3$ and $k=4$ to compute
$\frac{\gamma^{(2)}_{1}-\gamma^{(2)}_{3}}{\gamma^{(2)}_{4}-\gamma^{(2)}_{3}}=\frac{\pi_{1}-\pi_{3}}{\pi_{4}-\pi_{3}}.$
Likewise, for candidate $k=2$, we use its score distribution from judge $j=1$, i.e. $\gamma^{(1)}_{2}=\Pr(\hat{S}_{2}^{(1)}=1)$, relative to candidates $k=3$ and $k=4$ to compute
$\frac{\gamma^{(1)}_{2}-\gamma^{(1)}_{3}}{\gamma^{(1)}_{4}-\gamma^{(1)}_{3}}=\frac{\pi_{2}-\pi_{3}}{\pi_{4}-\pi_{3}}.$
Thus we have
$\frac{\gamma^{(1)}_{2}-\gamma^{(1)}_{3}}{\gamma^{(1)}_{4}-\gamma^{(1)}_{3}}=\frac{\pi_{1}-\pi_{3}}{\pi_{2}-\pi_{3}}.$
By ranking these shifted and scaled judge-assigned score distributions, we can recover the true ranking between all candidates. ∎
F.2 Necessary and Sufficient Conditions for 2-Level Systems
While Theorem 1 provides sufficient conditions for ranking identifiability under constancy assumptions, we can characterize the exact necessary and sufficient conditions for when rankings are preserved in 2-level systems with a single judge.
Lemma 1.
Let $(k_{1},k_{2},\ldots,k_{K})$ be a permutation such that the true scores satisfy $\Pr(S_{k_{1}}^{*}=2)>\Pr(S_{k_{2}}^{*}=2)>\cdots>\Pr(S_{k_{K}}^{*}=2)$. Then the following are equivalent:
  1. (i)
    The judge-assigned scores preserve this ordering: $\Pr(\hat{S}_{k_{1}}^{(j)}=2)>\Pr(\hat{S}_{k_{2}}^{(j)}=2)>\cdots>\Pr(\hat{S}_{k_{K}}^{(j)}=2)$
  2. (ii)
    For all $\ell\in\{1,\ldots,K-1\}$: $\Pr(\hat{S}_{k_{\ell}}^{(j)}=2)>\Pr(\hat{S}_{k_{\ell+1}}^{(j)}=2)$
Proof.
The forward direction (i $\Rightarrow$ ii) is immediate. For (ii $\Rightarrow$ i), assume $Pr(\hat{S}_{k_{\ell}}=2)>Pr(\hat{S}_{k_{\ell+1}}=2)$ for every $\ell=1,\dots,K-1$. Fix any indices $1\leq i<j\leq K$. By repeated transitivity,
$\hat{p}_{k_{i}}>\hat{p}_{k_{i+1}}>\cdots>\hat{p}_{k_{j}},$
hence $\hat{p}_{k_{i}}>\hat{p}_{k_{j}}$. Since this holds for all $i<j$, we obtain the strict chain $\hat{p}_{k_{1}}>\hat{p}_{k_{2}}>\cdots>\hat{p}_{k_{K}}$, proving (i). ∎
The key insight is that we only need to verify K1K-1 local inequalities between adjacent candidates rather than all $\binom{K}{2}$ pairwise comparisons. For condition (ii) to hold given the true ranking, we require:
Theorem 4 (Necessary and Sufficient Conditions).
Consider a 2-level scoring system with a single judge. True rankings are preserved by judge $j$ for all possible candidate configurations if and only if for each adjacent pair $(k_{\ell},k_{\ell+1})$ in the true ranking:
$\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}\quad\text{(cross-term dominance)}$ (9)
$\theta_{1,k_{\ell}}^{(1)}\geq\theta_{1,k_{\ell+1}}^{(1)},\text{ and if equal, then }\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell}}^{(1)}$ (10)
$\theta_{2,k_{\ell}}^{(1)}\geq\theta_{2,k_{\ell+1}}^{(1)},\text{ and if equal, then }\theta_{2,k_{\ell+1}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}$ (11)
where $\theta_{m,k}^{(1)}=\Pr(\hat{S}_{k}^{(1)}=2|S_{k}^{*}=m)$ represents the judge’s confusion matrix entries.
Proof.
Condition (ii) from Lemma 1 states that $\Pr(S_{k_{\ell}}^{*}=2)>\Pr(S_{k_{\ell+1}}^{*}=2)$ implies $\Pr(\hat{S}_{k_{\ell}}^{(1)}=2)>\Pr(\hat{S}_{k_{\ell+1}}^{(1)}=2)$. Using the notation $\pi_{k,m}=\Pr(S_{k}^{*}=m)$ and $\theta_{m,k}^{(1)}=\Pr(\hat{S}_{k}^{(1)}=2|S_{k}^{*}=m)$, this becomes:
$\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}\implies\gamma_{k_{\ell}}^{(1)}>\gamma_{k_{\ell+1}}^{(1)}$ (12)
where $\gamma_{k}^{(1)}=\theta_{1,k}^{(1)}(1-\pi_{k,2})+\theta_{2,k}^{(1)}\pi_{k,2}$ is the judge-assigned score probability.
For (12) to hold, we need the difference
$g(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2})$ $=\gamma_{k_{\ell}}^{(1)}-\gamma_{k_{\ell+1}}^{(1)}$
$=\theta_{1,k_{\ell}}^{(1)}(1-\pi_{k_{\ell},2})+\theta_{2,k_{\ell}}^{(1)}\pi_{k_{\ell},2}-\theta_{1,k_{\ell+1}}^{(1)}(1-\pi_{k_{\ell+1},2})-\theta_{2,k_{\ell+1}}^{(1)}\pi_{k_{\ell+1},2}$
$=(\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)})+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})\pi_{k_{\ell},2}+(\theta_{1,k_{\ell+1}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)})\pi_{k_{\ell+1},2}$ (13)
to be positive for all $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$ with $0\leq\pi_{k_{\ell+1},2}\leq\pi_{k_{\ell},2}\leq 1$.
Since $g(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2})$ is linear in both arguments, it is positive throughout the feasible region if and only if it is non-negative at all extreme points. The feasible region $\{(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2}):0\leq\pi_{k_{\ell+1},2}\leq\pi_{k_{\ell},2}\leq 1\}$ has three extreme points:
At $(0,0)$: $g(0,0)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}\geq 0$
At $(1,0)$: $g(1,0)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})=\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}>0$
At $(1,1)$: $g(1,1)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})+(\theta_{1,k_{\ell+1}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)})=\theta_{2,k_{\ell}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)}\geq 0$
Note that $g(1,0)$ must be strictly positive since this corresponds to the interior of the feasible region where $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$.
For strict inequality in (12) when $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$, we also need $g$ to be strictly positive when moving slightly away from the boundary points. This requires:
  • If $\theta_{1,k_{\ell}}^{(1)}=\theta_{1,k_{\ell+1}}^{(1)}$, then $\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell}}^{(1)}$ to ensure $g(\Delta,0)>0$ for small $\Delta>0$
  • If $\theta_{2,k_{\ell}}^{(1)}=\theta_{2,k_{\ell+1}}^{(1)}$, then $\theta_{2,k_{\ell+1}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}$ to ensure $g(1,1-\Delta)>0$ for small $\Delta>0$
These conditions correspond exactly to equations (9), (10), and (11). ∎
Remark 1.
These necessary and sufficient conditions reveal that without constancy assumptions, ranking preservation requires complex coordination between confusion matrices: higher-ranked candidates must have both larger false-positive rates (FPR) and true-positive rates (TPR) than lower-ranked candidates, with the higher rank’s TPR exceeding the lower rank’s FPR.
F.3 Proof for Theorem 2
We provide the proof for three levels. It can be easily extended to 3+ levels.
Proof for Theorem 2.
We first prove Theorem 2 for a single judge. It suffices to show that there exist two candidates for whom their relative rankings cannot be identified given only the judge-assigned score distribution.
Consider two candidates in the strict interior of the probability simplex, whose marginal judge-assigned score distributions are denoted $\gamma_{1}$ and $\gamma_{2}$. Let us first consider some judge whose vertices are denoted $\Theta=\left(\begin{matrix}\theta_{1}&\theta_{2}&\theta_{3}\end{matrix}\right)$, whose vertices are also strictly in the interior of the probability simplex and where $\Theta$ has full column rank. Then the distribution of the true scores (the barycentric coordinates) for candidate $k$ is given by $\pi_{k}=\Theta^{-1}\gamma_{k}$. The true scores of the two candidates are thus equal if
$\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)=0.$ (14)
Now it is possible that $\gamma_{1}$ and $\gamma_{2}$ were generated by a slightly different judge with triangle corners defined by $\Theta^{\prime}_{h}=\Theta+h\Delta$, where $h\in\mathbb{R}$ and $\Delta$ is any matrix such that the columns sum to zero. (As long as $h$ is sufficiently small, $\Theta^{\prime}_{h}$ is a valid set of judge vertices.) If this other judge were the true judge, then the difference of the true scores between the two candidates would be given by
$\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right).$ (15)
To prove nonidentifiability of the rankings, it thus suffices to show that for any $\Theta$, there exists $\gamma_{1},\gamma_{2},\Delta$ such that the score difference (15) is zero at $h=0$, i.e. (14) holds, and the derivative of the score difference (15) is nonzero, i.e.
$\nabla_{h}\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\propto\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}(\Theta^{-1}\Delta)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\neq 0.$ (16)
If this were to hold, then there exists some $h>0$ such that the true score rankings between two candidates with marginal score distributions $\gamma_{1}$ and $\gamma_{2}$ from a judge with vertices $\Theta-h\Delta$ would be the opposite if the judge instead had vertices $\Theta+h\Delta$.
To find such a $\gamma_{1},\gamma_{2}$, and $\Delta$, let $\bar{\pi}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})^{T}$ and $\pi_{1}=\bar{\pi}+\frac{\epsilon a}{2}$, $\pi_{2}=\bar{\pi}-\frac{\epsilon a}{2}$, where $a=(1,-2,1)^{T}$, and $\epsilon>0$ and small enough such that $\pi_{1}$ and $\pi_{2}$ belong to the probability simplex. Note that $\pi_{1}^{T}\mathbf{1}=\pi_{2}^{T}\mathbf{1}=\bar{\pi}^{T}\mathbf{1}=1$ since $a^{T}\mathbf{1}=0$.
[…]
$\Delta=(u-\bar{u}\mathbf{1})a^{\top}\in\mathbb{R}^{3\times 3}.$
Because $\mathbf{1}^{\top}(u-\bar{u}\mathbf{1})a^{T}=\vec{0}^{\top},$ we have constructed a $\Delta$ such that every column sums to 0. Moreover, because $\Theta^{-1}(\gamma_{1}-\gamma_{2})=\epsilon\alpha$, (16) simplifies to
$\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}(\Theta^{-1}\Delta)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)$ $=u^{T}\Delta\epsilon a=u^{T}(u-\bar{u}\mathbf{1})a^{T}a\epsilon$
$=\sum_{m=1}^{3}(u_{m}-\bar{u})^{2}\lVert a\rVert^{2}\epsilon$
$\geq 0.$
Now we prove that the inequality is in fact strict. We do this by contradiction. In particular, note that the equity holds if and only if $\vec{u}=(\Theta^{-1})^{T}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)=c\vec{1}$ for some $c\in\mathbb{R}$. Left-multiplying all elements in the equality by $\Theta^{\top}$, this is equivalent to assuming
$\Theta^{\top}\vec{u}=\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)=c\Theta^{\top}\vec{1}.$
However, because each column of $\Theta$ is a probability vector, this would imply that $(0,1,2)=c\vec{1}$, which would be a contradiction. This implies the score derivative is strictly nonzero.
Note that by adding more judges, the non-identifiability results remain the same. We can still find candidates and judges that satisfy the above conditions.
F.4 Proof for asymptotic consistency of independent model
As outlined in the main text, the Bayesian model assumes the probability model
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big].$ (17)
Here we prove that the candidate ranking that maximizes (17) still converges asymptotically to the true ranking as long as the true rank is identifiable, despite the probability model ignoring correlation between judge-assigned scores (and despite the fact that $\theta$ and $\pi$ are not identifiable). This can be viewed as an extension of traditional statistical results that the MLE for parameters that only depend on the first-moment is asymptotically consistent even when we do not correctly model the second-moments or higher. This result is stated more formally below. For simplicity, we present the result for one judge that satisfies Assumption 4. Extensions to multiple judges and to use Assumption 5 rather than 4 are straightforward.
Theorem 5.
Suppose Assumption 4 holds for one judge. The estimated ranking $\hat{\xi}_{n}$ using scores from the judge by maximizing (17) converges to the true ranking $\xi^{*}$ as $n\rightarrow\infty$ as long as:
  • the true ranking is identifiable from the marginals and
  • the true mean score for any two candidates are not exactly the same
Proof for Theorem 5.
Let us first consider a multinomial model of the judge-assigned scores where $(\gamma_{1},\cdots,\gamma_{K})$ are the paramerers of the judge-assigned score distributions, i.e.
$\max_{\gamma_{1},\cdots,\gamma_{K}}\prod_{i=1}^{n}\prod_{k=1}^{K}\Pr\left(\hat{S}_{i,k};\gamma_{k}\right).$ (18)
For notational convenience, let $\ell_{n}$ denote the log likelihood, i.e.
$\ell_{n}(\gamma_{1},\cdots,\gamma_{K})=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{K}\log\Pr\left(\hat{S}_{i,k};\gamma_{k}\right).$
Because the solution to (18) corresponds to the MLE for a multinomial model, and standard regularity conditions apply, the estimator is asymptotically consistent. Formally, this implies that the estimated parameters $\hat{\gamma}_{k,n}$ satisfy
$\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right)\rightarrow 1$ (19)
for any $a>0$.
Let’s analyze the optimization problem (17) optimized with respect to the parameters $(\theta,\delta,\xi)$ where $\xi$ is the ranking of the candidate, $\delta\in\mathbb{R}^{K,+}$, and $\pi_{k}=\sum_{i=1}^{K}\mathbbm{1}\{i<=\xi_{k}\}\delta_{i}$. Let a function $g$ denote the mapping from a parameter set $(\theta,\delta,\xi)$ describing judge and candidates to the set of marginal parameters $\gamma_{1},\cdots,\gamma_{K}$. Under Assumption 4, there is some true set of parameters $(\theta^{*},\delta^{*},\xi^{*})$ with corresponding marginal parameters $\gamma_{1}^{*},\cdots,\gamma_{K}^{*}$. Then because the set of models considered in (17) is a subset of that in (18), we have that
$\Pr\left(\ell_{n}(g(\theta,\delta,\xi))-\ell_{n}(g(\theta^{*},\delta^{*},\xi^{*}))<0\quad\forall\left\|g(\theta,\delta,\xi)-q(\theta^{*},\delta^{*},\xi^{*})\right\|\geq a\right)$ (20)
$\geq$ $\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right).$ (21)
Because (21) converges to one for any $a>0$, then (20) converges asymptotically to one as well for all $a>0$. Taking $a\rightarrow 0$, we have that
$\sup_{\theta_{n},\delta_{n},\xi_{n}}g(\hat{\theta}_{n},\delta_{n},\xi_{n})-(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})\rightarrow_{p}0.$ (22)
Furthermore, because we assumed that (i) the true ranking $\xi^{*}$ is identifiable from the marginals and (ii) no two candidates have the same exact true score, then the mapping from candidate marginals $(\gamma_{1},\cdots,\gamma_{K})$ to candidate rankings is unique and locally constant at $(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})$. Then by the continuous mapping theorem, we have that the estimated rankings $\hat{\xi}_{n}$ when maximizing (17) converges asymptotically to the true rankings $\xi^{*}$. ∎